# SciFIBench: Benchmarking Large Multimodal Models for Scientific Figure Interpretation

### v1

@@ Abstract
Large multimodal models (LMMs) have proven flexible and generalisable across many tasks and fields. Although they have strong potential to aid scientific research, their capabilities in this domain are not well characterised. A key aspect of scientific research is the ability to understand and interpret figures, which serve as a rich, compressed source of complex information. In this work, we present SciFIBench, a scientific figure interpretation benchmark. Our main benchmark consists of a 1000-question gold set of multiple-choice questions split between two tasks across 12 categories. The questions are curated from CS arXiv paper figures and captions, using adversarial filtering to find hard negatives and human verification for quality control. We evaluate 26 LMMs on SciFIBench, finding it to be a challenging benchmark. Finally, we investigate the alignment and reasoning faithfulness of the LMMs on augmented question sets from our benchmark. We release SciFIBench to encourage progress in this domain.111https://github.com/jonathan-roberts1/SciFIBench
Keywords:
AI4Science Benchmark LMMs Scientific Figures
@@ 1 Introduction
Lately, the rate of progress in the development of artificial intelligence (AI) has significantly increased. The emergence of foundation models [7], trained on large-scale broad data using extensive computational resources enabling generalisation across many downstream applications, has greatly expanded the range of possible domains and tasks in which machine intelligence can operate. Notable large language models (LLMs), such as GPT-4 [37], LLaMA [51], and PaLM [9], and subsequent large multimodal models (LMMs)222We use the term large multimodal model to refer to the family of models also known as multimodal large language models or large vision-language models. , for example, GPT-4V [38], Qwen [6], and Gemini [49], have proven to be flexible and generalisable across many tasks. In particular, their capabilities have been demonstrated in fields such as mathematics [14, 64, 60], medicine [27, 54, 12, 24, 60], and finance [55, 59], as well as writing code [8] and the geographic and geospatial domains [44, 45].
@@
We circumvent these issues by reframing the evaluation to a multiple-choice setting, using the figure captions as ground truth descriptions – see Fig. 1. Concretely, using 100k figure-caption pairs from arXiv papers, we construct a pool of multiple-choice questions for the two tasks shown in Fig. 1. Following other popular works [66], we adopt adversarial filtering when curating the negatives for each question to increase the difficulty. To further improve the quality, we utilise human verification on every question to ensure they are maximally answerable. We create SciFIBench (Scientific Figure Interpretation Benchmark) by sampling from this question pool with the following three objectives in mind: (1) Quality – we perform human verification on every question to ensure high-quality questions that are answerable. (2) Efficiency – we choose a small-scale set of questions, enabling streamlined evaluation and ensuring the benchmark can maximally be used by the community. (3) Robustness – we conduct careful analysis to verify SciFIBench offers a suitably robust evaluation. Our benchmark includes a gold dataset consisting of 1000 questions, a silver dataset of 10k questions, and a bronze dataset consisting of 174k questions (see Tab. 1).
We evaluate a suite of 26 open- and closed-source LMM baselines on SciFIBench and compare the performance to human and vision-language model (VLM) baselines. To overcome the challenges associated with post-processing the output of LMMs to extract a specific answer at scale, we leverage Gemini-Pro [49] to parse the output of all evaluated LMMs and extract the relevant multiple-choice letter answers, enabling automatic evaluation. Finally, we carry out preliminary experiments probing the alignment and faithfulness of the LMMs when answering questions in our benchmark. We hope our insights will encourage further research in this direction.
To conclude, our main contributions are as follows: (i) We curate SciFIBench to evaluate scientific figure interpretation. (ii) We benchmark 26 LMMs on SciFIBench and compare the performance to human and VLM baselines. (iii) We introduce an experimental setting probing the instruction-following abilities and faithfulness of reasoning of the LMMs. (iv) We release SciFIBench to drive progress in LMM scientific figure interpretation and understanding research.
We derive these key insights from our work:
@@
Closed-source LMMs outperform open-source models on our benchmark.
@@ 2 Related Work
2.1 Scientific Figure Interpretation
@@
Several approaches have been proposed to investigate the capacity of multimodal models to interpret scientific figures. These include question answering benchmarks such as ChartQA [34], PlotQA [35], and FigureQA [22], which ask complex reasoning questions about scientific figures. ACL-Fig [23] introduces the task of type classification for scientific figures. A large body of literature exists that evaluates the quality of generated captions for scientific figures. The progenitor for many subsequent works is SciCap [18], in which an image-captioning model is trained to generate high-quality captions. SciCap+ [61] builds this idea further and includes figure mention-paragraphs in addition to input figures. SciCap-Eval [19] investigates the usage of LLMs for ranking scientific figure captions. VisText [48] fine-tunes language models to generate captions for scientific charts, and FigCaps-HF [46] introduces a framework that initially learns a human feedback prediction model and incorporates this to optimise caption generation based on reader preference. The SciMMIR benchmark [56] characterises the abilities of vision-language models to understand scientific figures through retrieval experiments. More recently, a few works [38, 60] have conducted a qualitative analysis of LMM (specifically, GPT-4V [38]) performance on a small handful of scientific figures. We draw inspiration from these works, incorporating some of the methodological ideas. However, our work focuses on a quantitative evaluation of LMMs for the task of understanding scientific figures, which has yet to be reported. We also re-frame the task to a multiple-choice setting as this is more suitable for robust evaluation of LMMs.
2.2 LMM Benchmarks
@@
A number of benchmarks aimed at multimodal model evaluation have been developed in recent years. Prominent natural image benchmarks include LVLM-eHub [58], MMBench [32], MME [16], MM-Vet [63], and SEEDBench [29] and SEEDBench-2 [28], which both consist of multiple-choice questions across different domains and evaluation dimensions. A small-scale geographic and geospatial benchmark is introduced in [45]. LAMM [62] evaluates a variety of computer vision tasks on 2D natural images as well as 3D point clouds. Other benchmarks, such as HallusionBench [17], focus on the failure modes and hallucinations of the models. MathVista [33] introduces a mathematical reasoning in visual contexts metadataset, which includes scientific figures and charts. This benchmark contains similar image types to our work but has a different focus and uses different question types. The MMMU benchmark [65] includes multi-discipline college-level image-based problems and questions. Although limited to text, we take inspiration by the adversarial filtering approach taken in [66], in the curation of the multiple-choice questions in our work. Our work incorporates stylistic and methodological inspiration from these works but tackles a different image type with a different overall focus of scientific figure interpretation.
@@ 3 SciFIBench
3.1 Overview
Overall, SciFIBench is comprised of 188k questions, derived from figures and captions extracted from arXiv papers, curated into two multiple-choice tasks. As illustrated in Tab. 1, we split our dataset into 3 subsets, each with different purposes. We focus on the high-quality gold set in our evaluation and benchmarking, and reserve the noisier silver and bronze subsets for downstream hyperparameter tuning, few-shot examples, or fine-tuning.
| Split | # Questions | Human | Category | Difficulty | |
| Figure Caption | Caption Figure | curation? | balanced? | ||
| Bronze | 87k | 87k | ✗ | ✗ | Easy |
| Silver | 5k | 5k | ✗ | ✓ | Medium |
| Gold | 500 | 500 | ✓ | ✓ | Hard |
@@ 3.2 Tasks
SciFIBench consists of the following two core tasks related to the interpretation of scientific figures (illustrated in Fig. 1):
@@ 3.3 Curation methodology
We use the SciCap dataset [18] as our source of scientific figure-caption pairs. SciCap is a large-scale dataset consisting of figures and corresponding captions extracted from arXiv computer science papers between the years 2010 - 2020.
@@
Deduplication: We initially drop any captions (and corresponding figures) if they are duplicates of other captions.
2.
@@
Compute embeddings: We then use a variant of the CLIP model [42]333Specifically, we use the ViT-H-14-378-quickgelu model pretrained on the DFN-5B dataset[15] as it attains strong zero-shot performance across numerous datasets [20]. to compute embeddings for each figure-caption pair. After normalising, we concatenate the text and image embeddings to form joint embeddings, represented as vectors , where is equal to 2048.
3.
@@
Construct vector database: Using the Faiss library [13], we create a vector database of the joint embeddings.
4.
@@
Find nearest neighbours: For each embedding, we search for the nearest neighbours based on Euclidean distance. Concretely, given the set of database embeddings and a query embedding , we compute the nearest neighbours of as:
(1) -
5.
@@
Similarity filtering: To increase the likelihood the multiple-choice questions are answerable we remove very similar figure-caption pairs from our dataset (e.g., with minor formatting differences but no semantic difference) by dropping a sample () if its distance to the query embedding (i.e., ) falls below a threshold.
6.
@@
Question construction: For each selected figure-caption pair, we create multiple-choice questions using the nearest neighbours. For the Figure Caption task, we create target captions by randomly shuffling the true caption with the corresponding nearest neighbour captions. Similarly, for the Caption Figure task, we create the target figures by randomly shuffling the true figure with the corresponding nearest neighbour figures.
7.
@@
Categorisation: We categorise questions based on the arXiv category of the paper of the true figure-caption pair. Questions in the 10 most common categories are grouped individually while those in less common categories are labelled ‘other cs’; questions from cross-listed papers are labelled ‘cross-list’.
8.
@@
Difficulty filtering: For each question, we adopt the average distance score of the joint embeddings of the negatives to the true answer as a measure of difficulty. We sort the questions based on this difficulty.
9.
@@
Human verification: We sample the most difficult questions per category and perform human verification to select ‘answerable’ questions. We classify a question as answerable if it contains sufficient information for a domain expert to determine the single correct answer (i.e., questions with ambiguous choices; or references to context-dependent details, such as ‘Exp. 1’, ‘Config. 1’, etc. are disregarded). Minor text edits were made for a small subset of the questions to reduce ambiguity.
@@
Following these steps, we obtain a pool of high-quality answerable questions (600 per task). We evaluate GPT-4V [38] and Gemini-Pro Vision [49] on the pool and select questions that either model answers incorrectly. Then, we sample the remaining questions per category to create the gold subset of 500 questions per task. Due to some categories having few answerable questions in the pool, category balance was approached, but not exactly achieved in all cases – Fig. 2 illustrates the category representation per task. For example, although the ‘cs.AI’ category is in the top 10 most common, the pool of possible questions was dominated by figures/captions from a single paper; to avoid introducing bias, we only included 10 such questions per task. The silver subset was then constructed by taking the next 5000 most difficult questions per task, sampled across categories, without human checking, and the bronze set was then formed of all remaining questions. Example gold set questions for each task are shown in Fig. 3. These examples demonstrate the increased difficulty introduced by adversarially selecting the negative multiple-choice answers relative to randomly selected choices. Another observation is that determining the correct answer for some questions requires interpreting fine-grained detail, especially in the figures (correctly answering the Caption Figure example requires information in the figure legend).
3.4 Quality control
@@
A central motivation for focusing our evaluation on a small gold subset is to ensure high quality. As such, having manually checked each question, we conservatively estimate the noise level in the gold set to be at most a few percent. In these minority, questionable cases, we estimate there is a reasonable chance the questions can be answered with appropriate domain expertise. Based on spot checks, we estimate the noise level on the silver and bronze sets to be 20-25%. Across all sets, the noise rate is close to zero with randomly selected negatives.
Minor cosmetic errors, such as typos in captions or obscured axis labels, originating from the original SciCap data [18] were deliberately left unchanged when included in SciFIBench to increase question realism and difficulty.
3.5 Question difficulty
@@
Preliminary ablation studies on a randomly sampled set of questions showed that, for nearly all the LMMs evaluated, finding hard negatives for each question using nearest neighbours determined based on the joint-embedding similarity yields the most challenging questions, with lower accuracy scores on the joint-embedding neighbours than the single-modality neighbours. Fig. 4 outlines the difficulty distribution of questions in the gold and silver subsets, based on distance. The effect of adversarially selecting negatives compared to random selection can be seen in the disparity of the orange and blue distributions, with the adversarial negatives having a much lower mean distance and therefore higher difficulty. As expected from the curation process, the gold adversarial distribution is more challenging than the silver distribution.
@@ 4 Experiments
Through a variety of experiments, we evaluate the scientific figure interpretation abilities of a selection of closed- and open-source LMMs on our SciFIBench benchmark. We conduct our evaluations in a zero-shot setting and provide prompting and inference details in the following subsections. As it contains higher quality and more challenging questions, we focus our evaluation on the SciFIBench gold dataset.
@@ 4.1 Baselines
LMMs. In addition to the current frontier closed-source models (i) GPT-4V [38], (ii) GPT-4o [39] (iii) Gemini-Pro Vision [49], (iv) Gemini-Pro 1.5 [43], (v) Claude 3 {Opus, Sonnet and Haiku} [5], we evaluation the following open-source models on our benchmark: (vi) IDEFICS [26], (vii) Qwen-VL [6], (viii) Emu2 [47], (ix) TransCore-M [41], (x) InternLM-Composer [50], (xi) InternLM-Composer2 [50], (xii) CogVLM [52], (xiii) OmniLMM [40], (xiv) Yi [1], (xv) InstructBLIP [11], (xvi) Monkey [30], and (xvii) LLaVA-1.5 [31]. We use chat / instruction-tuned variants of each model (rather than base models) and compare the performance of multiple model sizes where available. Roughly half of these baselines can take interleaved text and images as input, and therefore be evaluated on the CaptionFigure task.
VLMs. As a point of comparison, we also evaluate strong VLM models on SciFIBench. Specifically, we evaluate a MetaCLIP [57] variant, the Google Multimodal Embedding Model444Google Vertex AI MM Embedding Model, 09/05/2024, and the CLIP model555https://huggingface.co/apple/DFN5B-CLIP-ViT-H-14-378 used to determine the nearest neighbour multiple-choice options.
Humans. Additionally, we evaluate a human baseline to gauge the relative performance difference between humans and LMMs. The humans (undergraduate and postgraduate students) were presented with the same prompt as the models.
While it is difficult to say with certainty if SciCap or arXiv data was included in the training sets of these models, there might be some leakage, as expected when using web images. However, given the scale of the training data, we do not expect this to impact our evaluation.

### v2

@@ Abstract
Large multimodal models (LMMs) have proven flexible and generalisable across many tasks and fields. Although they have strong potential to aid scientific research, their capabilities in this domain are not well characterised. A key aspect of scientific research is the ability to understand and interpret figures, which serve as a rich, compressed source of complex information. In this work, we present SciFIBench, a scientific figure interpretation benchmark consisting of 2000 questions split between two tasks across 8 categories. The questions are curated from arXiv paper figures and captions, using adversarial filtering to find hard negatives and human verification for quality control. We evaluate 28 LMMs on SciFIBench, finding it to be a challenging benchmark. Finally, we investigate the alignment and reasoning faithfulness of the LMMs on augmented question sets from our benchmark. We release SciFIBench to encourage progress in this domain: https://scifibench.github.io/.
@@ 1 Introduction
Lately, the rate of progress in the development of artificial intelligence (AI) has significantly increased. The emergence of foundation models [1], trained on large-scale broad data using extensive computational resources enabling generalisation across many downstream applications, has greatly expanded the range of possible domains and tasks in which machine intelligence can operate. Notable large language models (LLMs), such as GPT-4 [2], LLaMA [3], and PaLM [4], and subsequent large multimodal models (LMMs), for example, GPT-4V [5], Qwen [6], and Gemini [7], have proven to be flexible and generalisable across many tasks. In particular, their capabilities have been demonstrated in fields such as mathematics [8, 9, 10], medicine [11, 12, 13, 14, 10], and finance [15, 16], as well as writing code [17] and the geographic and geospatial domains [18, 19].
@@
We circumvent these issues by reframing the evaluation to a multiple-choice setting, using the figure captions as ground truth descriptions – see Fig. 1. Concretely, using figure-caption pairs from arXiv papers, we construct a pool of multiple-choice questions for the two tasks shown in Fig. 1. Following other popular works [30], we adopt adversarial filtering when curating the negatives for each question to increase the difficulty. To further improve the quality, we utilise human verification on every question to ensure they are maximally answerable. We create SciFIBench (Scientific Figure Interpretation Benchmark) by sampling from this question pool with the following three objectives in mind: (1) Quality – we perform human verification on every question to ensure high-quality questions that are answerable. (2) Efficiency – we choose a small-scale set of questions, enabling streamlined evaluation and ensuring the benchmark can maximally be used by the community. (3) Robustness – we conduct careful analysis to verify SciFIBench offers a suitably robust evaluation. Our benchmark consists of 2000 high-quality, challenging questions.
We evaluate a suite of 28 open- and closed-source LMM baselines on SciFIBench and compare the performance to human and vision-language model (VLM) baselines. To overcome the challenges associated with post-processing the output of LMMs to extract a specific answer at scale, we leverage Gemini-Pro [7] to parse the output of all evaluated LMMs and extract the relevant multiple-choice letter answers, enabling automatic evaluation. Finally, we carry out preliminary experiments probing the alignment and faithfulness of the LMMs when answering questions in our benchmark. We hope our insights will encourage further research in this direction.
To conclude, our main contributions are as follows: (i) We curate SciFIBench to evaluate scientific figure interpretation. (ii) We benchmark 28 LMMs on SciFIBench and compare the performance to human and VLM baselines. (iii) We introduce an experimental setting probing the instruction-following abilities and faithfulness of reasoning of the LMMs. (iv) We release SciFIBench to drive progress in LMM scientific figure interpretation and understanding research. We derive these key insights from our work:
@@ 2 Related Work
Scientific Figure Interpretation. Several approaches have been proposed to investigate the capacity of multimodal models to interpret scientific figures. These include question answering benchmarks such as ChartQA [33], PlotQA [34], and FigureQA [35], which ask complex reasoning questions about scientific figures. ACL-Fig [36] introduces the task of type classification for scientific figures. A large body of literature exists that evaluates the quality of generated captions for scientific figures.
@@
The progenitor for many subsequent works is SciCap [37], in which an image-captioning model is trained to generate high-quality captions. SciCap+ [38] builds this idea further and includes figure mention-paragraphs in addition to input figures. SciCap-Eval [39] investigates the usage of LLMs for ranking scientific figure captions. VisText [40] fine-tunes language models to generate captions for scientific charts, and FigCaps-HF [41] introduces a framework that initially learns a human feedback prediction model and incorporates this to optimise caption generation based on reader preference. The SciMMIR benchmark [24] characterises the abilities of vision-language models to understand scientific figures through retrieval experiments. More recently, a few works [5, 10] have conducted a qualitative analysis of LMM (specifically, GPT-4V [5]) performance on a small handful of scientific figures.
We draw inspiration from these works, incorporating some of the methodological ideas. However, our work focuses on a quantitative evaluation of LMMs for the task of understanding scientific figures, which has yet to be reported. We also re-frame the task to a multiple-choice setting as this is more suitable for robust evaluation of LMMs.
LMM Benchmarks. A number of benchmarks aimed at multimodal model evaluation have been developed in recent years. Prominent natural image benchmarks include LVLM-eHub [42], MMBench [26], MME [43], MM-Vet [44], and SEEDBench [28] and SEEDBench-2 [29], which both consist of multiple-choice questions across different domains and evaluation dimensions.
@@
A small-scale geographic and geospatial benchmark is introduced in [19]. LAMM [45] evaluates a variety of computer vision tasks on 2D natural images as well as 3D point clouds. Other benchmarks, such as HallusionBench [46], focus on the failure modes and hallucinations of the models. MathVista [27] introduces a mathematical reasoning in visual contexts metadataset, which includes scientific figures and charts. This benchmark contains similar image types to our work but has a different focus and uses different question types. The MMMU benchmark [25] includes multi-discipline college-level image-based problems and questions. Although limited to text, we take inspiration by the adversarial filtering approach taken in [30], in the curation of the multiple-choice questions in our work. Our work incorporates stylistic and methodological inspiration from these works but tackles a different image type with a different overall focus of scientific figure interpretation.
@@ 3 SciFIBench
SciFIBench is comprised of 2000 questions, derived from figures and captions extracted from arXiv papers, curated into two tasks and split into 2 subsets based on the data source (Tab. 1).
| Subset | # questions per task | Num. pairs | Num. | Source | Human | |
| FigureCaption | CaptionFigure | in pool | categories | data | verified? | |
| CS | 500 | 500 | 94k | 1 | [37] | ✓ |
| General | 500 | 500 | 102k | 7 | [47] | ✓ |
@@ 3.1 Tasks
SciFIBench consists of the following two tasks related to scientific figure interpretation (Fig. 1):
@@ 3.2 Curation methodology
We use the SciCap dataset [37] as our initial source of scientific figure-caption pairs. SciCap is a large-scale dataset consisting of figures and corresponding captions extracted from arXiv computer science (CS) papers between the years 2010-2020.
@@
1. Deduplication: We initially drop any captions (and corresponding figures) if they are duplicates.
2. Compute embeddings:
@@
We then use a variant of CLIP [48]111We use ViT-H-14-378-quickgelu [49] as it attains strong zero-shot performance on numerous datasets [50].. to compute embeddings for each figure-caption pair. After normalising, we concatenate the text and image embeddings to form joint embeddings, represented as vectors , where is equal to 2048.
@@
3. Construct vector database: Using Faiss [51], we create a vector database of the joint embeddings.
@@
4. Find nearest neighbours: For each embedding, we search for the nearest neighbours based on Euclidean distance. Concretely, given the set of database embeddings and a query embedding , we compute the nearest neighbours of as:
@@
5. Similarity filtering: To increase the likelihood the multiple-choice questions are answerable we remove very similar figure-caption pairs from our dataset (e.g., with minor formatting differences but no semantic difference) by dropping a sample () if its distance to the query embedding (i.e., ) falls below a threshold.
@@
6. Question construction: For each selected figure-caption pair, we create multiple-choice questions using the nearest neighbours. For the FigureCaption task, we create target captions by randomly shuffling the true caption with the corresponding nearest neighbour captions. Similarly, for the CaptionFigure task, we create the target figures by randomly shuffling the true figure with the corresponding nearest neighbour figures.
@@
7. Categorisation: We categorise questions based on the arXiv category of the true figure-caption pair. Questions in the 10 most common categories are grouped individually while those in less common categories are labelled ‘other cs’; questions from cross-listed papers are labelled ‘cross-list’.
@@
8. Difficulty filtering: We adopt the average distance of the joint embeddings of the negatives to the true answer as a measure of question difficulty. We sort the questions based on this difficulty.
@@
9. Human verification: We sample the most difficult questions per category and perform human verification to select ‘answerable’ questions. We classify a question as answerable if it contains sufficient information for a domain expert to determine the single correct answer (i.e., questions with ambiguous choices; or references to context-dependent details, such as ‘Exp. 1’, ‘Config. 1’, etc. are disregarded). Minor text edits were made for a small subset of the questions to reduce ambiguity.
category representation.
@@
Following these steps, we obtain a pool of high-quality questions. We evaluate GPT-4V [5] and Gemini-Pro Vision [7] on the pool and select questions that either model answers incorrectly and sample the remaining questions per category to create our curated ‘CS’ question set. As some categories had few answerable questions in the pool, category balance was approached, but not achieved in all cases – Fig. 5 illustrates the category representation. For example, the pool of possible ‘cs.AI’ category questions was dominated by figures/captions from a single paper; to avoid introducing bias, we only included 10 such questions per task. For analysis, a noisier subset was then constructed by taking the next 5000 most difficult questions per task, sampled across categories, without human checking. Example questions from the curated set are shown in Fig. 5.
To expand the diversity of the questions in our benchmark, we utilised the ArXivCap dataset [47], which contains figure-caption pairs from papers covering 32 arXiv domains up to 2023. Due to its larger scale, we initially randomly downselected 25% of the data and removed all pairs from CS papers. We then repeated the curation steps outlined above, to create a pool of answerable questions for each task. We then carried out a category-balanced downsampling of questions Gemini 1.5 Pro [32] and GPT-4o [31] answered incorrectly to reach a final curated ‘General’ set of 500 questions per task. In Fig.2 we show figure and caption statistics for questions across both subsets of SciFIBench.
@@
Quality control. We focus our evaluation on a small set of questions to ensure high quality. Having manually checked each question, we conservatively estimate the noise level in the curated set – i.e. included in SciFIBench – to be at most a few percent. In these minority, questionable cases, we estimate there is a reasonable chance the questions can be answered with appropriate domain expertise. Based on spot checks, we estimate the noise level on the noisy, uncurated questions to be 20-25%. Minor cosmetic errors, such as typos in captions or obscured axis labels, originating from the original data were deliberately left unchanged when included in SciFIBench to increase realism and difficulty.
@@
Question difficulty. Preliminary ablation studies on a random set of questions showed that, for nearly all the LMMs evaluated, selecting hard negatives using nearest neighbours determined by joint-embedding similarity yields the most challenging questions, with lower accuracy scores than the single-modality neighbours. Fig. 5 outlines a comparison of the difficulty distribution of the included curated questions and uncurated noisy questions, based on distance. The effect of adversarial compared to random negatives selection can be seen in the disparity of the orange and blue distributions, with the adversarial negatives having a much lower mean distance and therefore higher difficulty. As expected from the curation process, the curated adversarial distribution is more challenging than the noisy distribution.
@@ 4 Experiments
Through a variety of experiments, we evaluate the scientific figure interpretation abilities of a selection of LMMs on our SciFIBench benchmark and conduct a detailed analysis of their performance.
@@ 4.1 Baselines
LMMs. We evaluate the following closed-source models: GPT-4 {V, Turbo, o} [5, 52, 31] Gemini-Pro Vision [7], Gemini 1.5 {Flash, Pro} [32], and Claude 3 {Opus, Sonnet and Haiku} [53]. We also evaluate the following open-source models: IDEFICS [54], Qwen-VL [6], Emu2 [55], TransCore-M [56], InternLM-Composer 1,2 [57], CogVLM [58], OmniLMM [59], Yi [60], InstructBLIP [61], Monkey [62], and LLaVA-1.5 [63]. We use chat / instruction-tuned variants of each model (rather than base models) and compare the performance of multiple model sizes where available. Roughly half of these baselines can take interleaved text and images as input, and therefore be evaluated on the CaptionFigure task. We also consider a text-only baseline in which we provide the LMM with the output from an OCR model [64] rather than images. VLMs. As a point of comparison, we evaluate strong VLM models on SciFIBench. Specifically, we evaluate a MetaCLIP [65] variant, the Google Multimodal Embedding Model [66], and the CLIP model [49] used to determine the nearest neighbour multiple-choice options. Humans. Additionally, we evaluate a human baseline to gauge the relative performance difference between humans and LMMs. The humans (undergraduate and postgraduate students) were presented with the same prompt as the models.
While it is difficult to say with certainty if arXiv data was included in the training sets of these models, there might be some leakage, as expected when using web images. However, given the scale of the training data, we do not expect this to impact our evaluation.